OpenStudy (anonymous):

A gold miner is mining for gold in Texas, California and Nevada. It is predicted there is a 52% chance of finding gold in Texas, a 31% chance in California and a 20% chance in Nevada. What is the probability that: A) Gold is found in California, but not Nevada? B) Gold is found in Nevada, but not Texas? C) Gold is not found in Texas, California nor Nevada? D) Gold is found in at least one of these states? Round to three decimal places. Write a decimal such as 0.25 as just .25.

OpenStudy (anonymous):

no one wants to do this one i see. let me draw a little venn diagram to make it easier

OpenStudy (zarkon):

Somewhere is should be stated that you are assuming Independence.

OpenStudy (anonymous):

no it doesn't state that

OpenStudy (zarkon):

then you can't do the problem

OpenStudy (anonymous):

it is certainly assuming independence. let's do A) \[P(C)=.31, P(N)=.2\] \[P(C\cap N^c)=P(C)-P(N\cap C)\] and \[P(N\cap C)=P(C)\times P(N)\] assuming independence. first off is this notation clear?

OpenStudy (anonymous):

yeah makes sense

OpenStudy (zarkon):

one can also go directly to this \[P(C\cap N^{c})=P(C)\cdot P(N^{c})\] by independence

OpenStudy (anonymous):

@satellite73 we understand your formulas we just dont know how to go about solving them.

OpenStudy (anonymous):

i am trying to find a nice venn diagram template. i am sure i had one but i cannot seem to find it. in any case we can compute \[P(C\cap N^c)=.31-.2\times .31=.248\]

OpenStudy (anonymous):

is that a

OpenStudy (anonymous):

similarly for \[P(N\cap T^c)=.2-.2\times .52=.098\]

OpenStudy (anonymous):

those are a and b respectively

OpenStudy (anonymous):

so A is .248 and B is .098

OpenStudy (zarkon):

.096

OpenStudy (anonymous):

for wat?

OpenStudy (zarkon):

.2-.2*.52=.096

OpenStudy (anonymous):

zarkon is right. it should be \[.2-.2\times .52=.096\]

OpenStudy (anonymous):

ok thanks and wats C and D?

OpenStudy (anonymous):

c) is a little tricky, unless i am missing something simple. \[P(T\cup N\cup C)=P(T)+P(N)+P(C)\] \[-P(T\cap N)-P(T\cap C) -P(N\cap C)+P(N \cap T\cap C)\]

OpenStudy (anonymous):

@zarcon if i am making too much of this jump in and correct it please

OpenStudy (zarkon):

\[P(T^{c}\cap C^{c} \cap N^{c})\] \[=P(T^{c})P(C^{c})P(N^{c})=.48\cdot .69\cdot.8=.26496\] again by independence.

OpenStudy (anonymous):

buy i think you have to compute \[.52+.31+.20-.52\times .31-.52\times .2-.31\times .2+.52\times .31\times .20\]

OpenStudy (anonymous):

ah yes zarcon way much easier sorry

OpenStudy (anonymous):

and D is \[1-c\]

OpenStudy (zarkon):

yes ;)

OpenStudy (zarkon):

I have a Ph.D. in mathematical statistics...I know what I am doing :)

OpenStudy (anonymous):

So the final answers are- A) .248, B) .096, C) .265 and D) .735

OpenStudy (anonymous):

@Zarkon are those answers right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you! stay around we got one more question for you

OpenStudy (anonymous):

Four Olympic skaters are attempting to win a medal. The skaters success rates for skaters A, B, C and D are: 80%, 83%, 52% and 86% respectively. Find the probability that: A) Skater A and D are successful, but C and B are not. B) Skater B and C are successful, but A and D are not. C) None of the skaters win a medal. D) At least one of the four skater wins a medal. Round to four decimal places. Write a decimal such as 0.25 as just .25.

OpenStudy (zarkon):

Again, assuming independence (which seems really weird in this case) \[P(A\text{ and }D\text{ and }C^{c}\text{ and }B^{c})\] \[P(A)P(D)P(C^{c})P(B^{c})\]

OpenStudy (zarkon):

these problems are done just like the last problem you posted

OpenStudy (anonymous):

well we still need help, please

OpenStudy (zarkon):

tell me what \[P(A)=\] \[P(D)=\] \[P(C^{c})=\] \[P(B^{c})=\]

OpenStudy (anonymous):

idk

OpenStudy (zarkon):

"The skaters success rates for skaters A, B, C and D are: 80%, 83%, 52% and 86% respectively. "

OpenStudy (zarkon):

are you telling me that you don't know at least the first two because they are written in the text I quoted from you.

OpenStudy (anonymous):

no we do but we thought you already knew

OpenStudy (anonymous):

we got A its .056 right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so wat are b, c and d

OpenStudy (zarkon):

last one I'm giving... b) .0120848

OpenStudy (anonymous):

but i need help

OpenStudy (zarkon):

tell me what you have and I'll tell you if it is correct. Again, these problems are done exactly like the problem you posted before

OpenStudy (anonymous):

the things is we don't know how to do it, we didn't quite understand your explanation last time

OpenStudy (zarkon):

for (c) you don't want anyone to get a metal so you need find the probability of the complement that do win the metal \[P(Z^{c})=1-P(Z)\] and multiply all four of them together

OpenStudy (anonymous):

what is PZ

OpenStudy (anonymous):

uhhh i just dont understand I suck at math

OpenStudy (zarkon):

um..ok.. take each of the probabilities given and subtract them from 1 them multiply those 4 numbers together

OpenStudy (anonymous):

.0022848

OpenStudy (zarkon):

yes...that is the answer, though you really have no idea why.

OpenStudy (zarkon):

the answer to part D is just 1 minus the answer to part C

OpenStudy (anonymous):

hahah no unfortunatly but .9977152

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

it was wrong but thanks anyway

OpenStudy (zarkon):

under the assumption of independence it is not wrong